A bat traveling at 10 m/s for 10 seconds has traveled how much distance? 


1
Expert's answer
2020-08-31T13:12:42-0400

The bat is travelling at constant velocity "v=10" m/s. We know that

"v = \frac{ds}{dt} \Rightarrow s = v \cdot t"

So, "s = 10 \; \frac{m}{s} \cdot 10 \; s = 100 \; m."

Answer: 100 m.